please help solve
a box with no top is to be constructed from a piece of cardboard whose length measures 6in. more than its width. The box is to be formed by cutting squares that measure 2in. on each side from the four corners, and then folding up the sides. If the volume of the box will be 110in.^3 what are the dimensions of the piece of cardboard?

Answer by venugopalramana(3286)

please help solve
a box with no top is to be constructed from a piece of cardboard whose length (L SAY)measures 6in. more than its width(W SAY)
L=W+6
. The box is to be formed by cutting squares that measure 2in. on each side from the four corners,
BOX LENGTH = L-2-2=L-4 =W+6-4=W+2
BOX WIDTH = W-2-2=W-4
BOX HEIGHT=2
and then folding up the sides. If the volume of the box will be 110in.^3 what are the dimensions of the piece of cardboard?
VOLUME OF BOX = LWH =(W+2)(W-4)2=110
W^2-2W-8=55
W^2-9W+7W-63=0
W(W-9)+7(W-9)=0
(W-9)(W+7)=0
W=9
HENCE L=W+6=15
CARD BOARD SIZE = 15 X 9 INCHES